(20 pts) A certain country only produces Graham Crackers and Chocolate Bars. Its production for the previous three years is below . Graham Crackers Chocolate Bars Year 2013 2014$2 201532 Price Quantity 50 75 100 Price $3 $4 $5 Quantity 100 130 200 $1 $4 Calculate the Nominal GDP and Real GDP for each year, using 2013 as the base year.
Answer #1

43 Nominal GDP2013 = Price2013*Qty2013 = $1*50 (crackers)+ $3 *100 (chocolate bars) = 50+300 = $350
nominal GDP2014 = Price2014*Qty2014 = $2 * 75 + $4 * 130 = $670
nominal GDP2015 = Price2015*Qty2015 = $4*100 + $5*200 = $1400

For the base year, nominal GDP is equal to real GDP. This means for 2013, Real GDP2013 will be $350.
Real GDP2014 = Price2013*Qty2014 = $1*75 + $3*130 = $465
  Real GDP2015 = Price2013*Qty2014 = $1*100 + $3*200 = $700